Two billiard balls of equal mass undergo a perfectly elastic head-on collision. If one ball's initial speed was 1.60 m/s , and the other's was 3.60 m/s in the opposite direction, what will be their speeds after the collision?
Answer #1

by conservation of momentum

initial momentum = final momentum

m * 1.6 + m * 3.6 = m * v1 * m * v2

1.6 + 3.6 = v1 * v2

by conservation of energy

initial energy = final energy

0.5 * m * 1.6^2 + 0.5 * m * 3.6^2 = 0.5 * m * v1^2 + 0.5 * m * v2^2

1.6^2 + 3.6^2 = v1^2 + v2^2

on solving we'll get

speed of billiard balls after collision

v1 = 1.414 m/s

v2 = 3.677 m/s